The diagonal of a rectangle is 10 inches and the area is 45 square inches. Find the dimensions of the triangle, correct to one decimal place.
Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

let l be length and w be width of the rectangle, then
:
1) l^2 + w^2 = 10^2
2) l * w = 45
:
we have two equations in two unknowns
:
solve 2 for l and substitute in 1
l = 45/w
(45/w)^2 + w^2 = 100
:
(2025/w^2) + w^2 = 100
:
multiply both sides of = by w^2
:
2025 + w^4 = 100w^2
:
w^4 -100w^2 + 2045 = 0
:
let u = w^2
:
u^2 -100u + 2025 = 0
:
use quadratic equation
:
u = (100 + square root( (-100)^2 - 4(1)(2025) ) ) / 2 = 71.8
u = (100 - square root( (-100)^2 - 4(1)(2025) ) ) / 2 = 28.2
:
since u = w^2, we have
:
w^2 = 71.8 and w = 8.5
w^2 = 28.2 and w = 5.3
:
we have 2 solutions for l as well
:
l = 45/8.5 = 5.3
l = 45/5.3 = 8.5
:
Our 2 solutions are
**********************
w = 8.5 and l = 5.3
w = 5.3 and l = 8.5
*********************
:
note that
8.5 * 5.3 = 45.05
8.5^2 + 5.3^2 = 100.34 and square root of 100.34 = 10.02
:

Answer by ikleyn(52787)

.
The data relates to a rectangle.

The question relates to an unknown triangle.

What it is this about? To vast our time?